Paris Manufacturing that produces computer hard drives has had problems with hard drive defects in the past. The relevant information on their current inspection system follows. Current Manual Inspection System Annual fixed cost = $32,000 Inspection variable cost per unit = $11.50 A marketing representative of NEW-SPEC, a firm that specializes in providing manual inspection processes for other firms, approached Paris Manufacturing and offered to inspect parts for $25.00 each with no fixed cost. It assured Paris Manufacturing that the accuracy and quality of its manual inspections would equal that of the automated inspection system. Demand for the upcoming year is forecast to be 10,000 units. Should the manufacturer accept the offer? This problem includes break-even calculations. For a review of how to calculate break-even analysis, please refer to Chapter 6 – Equations 6.1, 6.2, and 6.3. Round your answer for the breakeven volume to the nearest whole number and round your answer for the amount of saving/loss to the nearest dollar. what is the break even volume per hard drives?
Answer #1

Given: Annual Demand = 10,000 units

For Manual Inspection System

Annual Fixed cost = $32,000

Inspection variable cost per unit = $11.50

Total cost for 10,000 units = Annual Fixed cost + Inspection variable cost per unit x Annual Demand = 32000+11.50 x 10000 = $147000

For Automated inspection system

Inspection variable cost per unit = $25

Total cost for 10,000 units = Annual Fixed cost + Inspection variable cost per unit x Annual Demand = 0 + 25 x 10000 = $250000

Since total cost for automated inspection system is more compared to manual system, hence the manufacturer should not accept the offer

For Break even volume

Let the annual demand be D

Hence, Total cost for 10,000 units for manual inspection system = Total cost for 10,000 units for automated inspection system

32000 + 11.50 x D = 25 x D

Hence, 13.5 x D = 32000

D = 2370.37 hard drives

If the Annual demand of hard drives is less than 2370, then the manufacturer should go for automated inspection system.

Hence, Savings for manufacturer by using manual inspection system = Total cost for 10,000 units for manual inspection system - Total cost for 10,000 units for automated inspection system = 250,000 - 147,000 = $103,000